OpenStudy (anonymous):

Evaluate the integral. x sin 3x dx from 0 to 1

OpenStudy (anonymous):

No. It's x sin (3x) dx

OpenStudy (anonymous):

sub u=3x

OpenStudy (australopithecus):

I'm pretty sure this can be solved by integration by parts where, f(x) = x g'(x) = sin(3x) watch this video https://www.youtube.com/watch?v=dqaDSlYdRcs maybe it can explain it bus basically \[\int\limits_{}^{}g'(x)f(x)dx = g(x)f(x) - \int\limits_{}^{}f'(x)g(x)dx\]

OpenStudy (anonymous):